A. Find the values of a and b that would make the quadrilateral a parallelogram. (Remember the properties of parallelograms.)
B. Make sure to show your setup and work to solve for both a and b in a neat, orderly fashion.
Top...6b-3
Leftside 2a+6
Rightside...3a-10
Bottom....5a+1

Answer by stanbon(75887)

Top...6b-3
Leftside 2a+6
Rightside...3a-10
Bottom....5a+1
-----
Equations:
6b-3 = 5a+1
2a+6 = 3a-10
-----------
a = 16
-----
6b - 5a = 4
---
6b -80 = 4
6b = 84
b = 14
-----------
Cheers,
Stan H.
------------